Two electric charges , originally 80mm apart , are brought closer until the force between them becomes greater by a factor of 9. How far apart are they now ?
1
Expert's answer
2018-01-28T09:30:07-0500
The Coulomb's law states
F=k (q_1 q_2)/r^2 .
Because F^'=9F so r^'=r/3.
Finally
r^'=(80 mm)/3=26.7 mm